OpenStudy (anonymous):

hey does anyone know if the following is convergent or divergent?? a(n)=( n/cos(n)) as n goes from 1 to infinity plz help:)

OpenStudy (anonymous):

If it's convergent \[\lim _{n \to \infty} {n \over \cos n} = L\] Since the limit of n to infinity is either infinity or minus infinity in this case, it does not converge to some limit L and is divergent

OpenStudy (anonymous):

thanks, that's really helpfull:)

OpenStudy (anonymous):

also, does anyone know if the following are true or false: (i) every sequence of real numbers is either bounded above or bounded below or both. (ii) some convergent sequences are not bounded. :)

OpenStudy (anonymous):

what is the fourth term of the sequence {max({0, (-2)^n})} as n goes from 1 to infinity and i also need to know the sixth term of the sequence, {min({0, (-2)^ (n+1) })} as n goes from 1 to infinity thanks:)